# Convergence of empirical Gromov-Wasserstein distance

### v1

@@ 4. Proofs
We start with verifying that one can choose points $\{x_{1},\ldots,x_{k}\}$ such that $\mu$ has mean zero. For a given integer $k$, consider a maximal set of points $\{x_{1}^{\prime},\ldots,x_{1}^{\prime}\}$ inside $B_{\mathbb{R}^{d}}(0,1/3)$ such that $\|x_{i}^{\prime}-x_{j}^{\prime}\|>\gamma_{d}k^{-1/d}$ for all $i\neq j$, where $\gamma_{d}$ is a small positive constant that depends only on $d$. Let $x_{i}:=x_{i}^{\prime}-\bar{x}^{\prime}$ with $\bar{x}^{\prime}=k^{-1}\sum_{i=1}^{k}x_{i}^{\prime}$, then $\{x_{1},\ldots,x_{k}\}\subset B_{\mathbb{R}^{d}}(0,2/3)$, $\left\lVert x_{i}-x_{j}\right\rVert>\gamma_{d}k^{-1/d}$ for all $i\neq j$, and the distribution $\mu=k^{-1}\sum_{i=1}^{k}\delta_{x_{i}}$ has mean zero.
@@
$S_{\delta}:=\{x\in\mathbb{R}^{d}:|v^{\top}x|\leq\delta\}$
@@
$\begin{split}D(\mu,\nu)&=8-8+32\epsilon-32\epsilon^{2}\\ &=32\epsilon(1-\epsilon),\end{split}$
@@
$\left\lVert x-y\right\rVert\geq\frac{1}{3}ck^{-1/d}\mathbbm{1}\{x\neq f(y)\}$
@@
$\begin{split}\int\left\lVert x-y\right\rVert\,d\pi(x,y)&\geq\frac{1}{3}ck^{-1/d}\mathbb{P}_{\pi}(X\neq f(Y))\\ &\geq\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(F_{\#}\mathfrak{q},(f\circ U\circ F)_{\#}\mathfrak{u})\\ &=\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\\ &=\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u}),\end{split}$
@@
[…]
@@
$\begin{split}&\sup_{(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))}\mathbb{P}\big{(}|\hat{D}_{n,m}^{1/2}-D^{1/2}(\mu,\nu)|\geq\Delta_{d}\big{)}\\ &\quad\geq\frac{1}{2}\left(\sup_{\mathfrak{q}\in\mathcal{D}_{k}^{+}}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}A)\big{]}+\sup_{\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}A)\big{]}\right)\\ &\quad\geq\frac{1}{2}\left(\sup_{{\mathfrak{q}\in\mathcal{D}_{k}^{+}}}\mathbb{P}_{\mathfrak{q}}(\psi=1)+\sup_{\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}}\mathbb{P}_{\mathfrak{q}}(\psi=0)\right)-.1\,.\end{split}$
@@ Appendix A Technical tools
Let $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ be a nonnegative lower semicontinuous function such that there exist nonnegative measurable functions $c_{\mathcal{X}}:\mathcal{X}\to\mathbb{R}_{+}$ and $c_{\mathcal{Y}}:\mathcal{Y}\to\mathbb{R}_{+}$ with $c\leq c_{\mathcal{X}}\oplus c_{\mathcal{Y}}$. Set $\mathcal{B}_{\mathcal{X}}(r)=c_{\mathcal{X}}^{-1}([0,r])$ and $\mathcal{B}_{\mathcal{Y}}(r)=c_{\mathcal{Y}}^{-1}([0,r])$ for $r>0$. Suppose that there exist constants $\kappa>0,\alpha\in(0,1/2]$ and $\delta>0$ such that
$\sup_{(\mu,\nu)\in\mathcal{P}(\mathcal{B}_{\mathcal{X}}(r))\times\mathcal{P}(\mathcal{B}_{\mathcal{Y}}(r))}\mathbb{E}\left[\big{|}T_{c}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c}(\mu,\nu)\big{|}\right]\leq\kappa r(n\wedge m)^{-\alpha}\big{(}\log_{+}(n\vee m)\big{)}^{\delta}$ (22)
@@
$\sup_{(\mu,\nu)\in\mathcal{Q}_{\epsilon,M}}\mathbb{E}\left[\big{|}T_{c}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c}(\mu,\nu)\big{|}\right]\lesssim(n\wedge m)^{-\alpha}\big{(}\log_{+}(n\vee m)\big{)}^{\delta}.$
The hidden constant depends only on $\kappa,\alpha,\epsilon$ and $M$.
@@
We follow the notation used in the proof of Theorem 1.1 in [SH25]. In our case, we may take $s=2$, which yields $\beta=\gamma=1/2$. As such, $\alpha\leq\beta$, so the conclusion follows from Comment 1 after the proof of Theorem 1.1 in [SH25]. ∎

### v2

@@ 4. Proofs
We start with verifying that one can choose points $\{x_{1},\ldots,x_{k}\}$ such that $\mu$ has mean zero. For a given integer $k$, consider a maximal set of points $\{x_{1}^{\prime},\ldots,x_{k}^{\prime}\}$ inside $B_{\mathbb{R}^{d}}(0,1/3)$ such that $\|x_{i}^{\prime}-x_{j}^{\prime}\|>\gamma_{d}k^{-1/d}$ for all $i\neq j$, where $\gamma_{d}$ is a small positive constant that depends only on $d$. If we let $x_{i}:=x_{i}^{\prime}-\bar{x}^{\prime}$ with $\bar{x}^{\prime}=k^{-1}\sum_{i=1}^{k}x_{i}^{\prime}$, then $\{x_{1},\ldots,x_{k}\}\subset B_{\mathbb{R}^{d}}(0,2/3)$, $\left\lVert x_{i}-x_{j}\right\rVert>\gamma_{d}k^{-1/d}$ for all $i\neq j$, and the distribution $\mu=k^{-1}\sum_{i=1}^{k}\delta_{x_{i}}$ has mean zero.
@@
$S_{\delta}:=\big{\{}x\in\mathbb{R}^{d}:|v^{\top}x|\leq\delta\big{\}}$
@@
$D(\mu,\nu)=8-8+32\epsilon-32\epsilon^{2}=32\epsilon(1-\epsilon),$
@@
$\|x-y\|\geq\frac{1}{3}ck^{-1/d}\mathbbm{1}_{\{x\neq f(y)\}}$
@@
$\begin{split}\int\|x-y\|\,d\pi(x,y)&\geq\frac{1}{3}ck^{-1/d}\mathbb{P}_{\pi}(X\neq f(Y))\\ &\geq\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(F_{\#}\mathfrak{q},(f\circ U\circ F)_{\#}\mathfrak{u})\\ &=\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\\ &=\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u}),\end{split}$
@@
$\begin{split}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}(A)\big{]}&\geq\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}\{\hat{D}_{n,m}^{1/2}>2\Delta_{d}\}\cap\{D^{1/2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\leq\Delta_{d}\}\big{)}\big{]}\\ &\geq\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}\hat{D}_{n,m}^{1/2}>2\Delta_{d}\big{)}\big{]}-\mathbb{P}\big{(}D^{1/2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})>\Delta_{d}\big{)}\\ &\geq\mathbb{P}_{\mathfrak{q}}(\psi=0)-0.1,\end{split}$
@@
$\begin{split}&\sup_{(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))}\mathbb{P}\big{(}|\hat{D}_{n,m}^{1/2}-D^{1/2}(\mu,\nu)|\geq\Delta_{d}\big{)}\\ &\quad\geq\frac{1}{2}\left(\sup_{\mathfrak{q}\in\mathcal{D}_{k}^{+}}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}A)\big{]}+\sup_{\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}A)\big{]}\right)\\ &\quad\geq\frac{1}{2}\left(\sup_{{\mathfrak{q}\in\mathcal{D}_{k}^{+}}}\mathbb{P}_{\mathfrak{q}}(\psi=1)+\sup_{\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}}\mathbb{P}_{\mathfrak{q}}(\psi=0)\right)-0.1.\end{split}$
@@
4.6. Proofs for Section 3.3
Proof of Lemma 3.1.
Set $a=\sum_{i=1}^{n}\mathsf{c}_{i}^{2}$ and $b=\sum_{i=1}^{N}\mathsf{c}_{i}$ for the notational convenience. Proposition 2 in [Com24] shows that
$\mathbb{P}\Big{(}\mathfrak{f}(Z)\geq\mathbb{E}[\mathfrak{f}(Z)\mid Z\in\mathcal{W}]+t\sqrt{a/2}+\mathfrak{p}b\Big{)}\leq\mathfrak{p}+e^{-t^{2}}.$
Observe that
$\mathbb{E}[\mathfrak{f}(Z)\mid Z\in\mathcal{W}]=\frac{1}{1-\mathfrak{p}}\mathbb{E}\big{[}\mathfrak{f}(Z)\mathbbm{1}_{\{Z\in\mathcal{W}\}}\big{]}\\ \leq\frac{1}{1-\mathfrak{p}}\mathbb{E}[\mathfrak{f}(Z)]\\ \leq 2\mathbb{E}[\mathfrak{f}(Z)],$
where the last inequality used the assumption that $\mathfrak{p}\leq\frac{1}{2}$. ∎
Proof of Theorem 3.3.
As before, we assume without loss of generality that $\mu$ and $\nu$ have mean zero. Recall that $\tilde{\mu}_{n}$ and $\tilde{\nu}_{n}$ are the centered versions of $\hat{\mu}_{n}$ and $\hat{\nu}_{n}$, respectively. Observe that
$\begin{split}\Delta_{n}&\leq|S_{1}(\tilde{\mu}_{n},\tilde{\nu}_{n})-S_{1}(\mu,\nu)|+|S_{2}(\tilde{\mu}_{n},\tilde{\nu}_{n})-S_{2}(\mu,\nu)|\\ &=:\mathfrak{f}(Z)+\mathfrak{g}(Z),\end{split}$
where $Z=(Z_{1},\dots,Z_{N}):=(X_{1},\dots,X_{n},Y_{1},\dots,Y_{n})\in\mathcal{X}^{n}\times\mathcal{Y}^{n}=:\mathcal{Z}$ with $N:=2n$. For $r\geq 1$, set
$\mathcal{W}_{r}:=\big{\{}z=(z_{1},\dots,z_{N})\in\mathcal{Z}:\max_{1\leq i\leq N}\|z_{i}\|\leq r\big{\}}$
and $\mathfrak{p}_{r}:=\mathbb{P}(Z\notin\mathcal{W}_{r})$. For any $Z,Z^{\prime}\in\mathcal{W}_{r}$ that differ only at the $i$-th coordinate for some $i\in\{1,\dots,N\}$, we shall show that
$|\mathfrak{f}(Z)-\mathfrak{f}(Z^{\prime})|\vee|\mathfrak{g}(Z)-\mathfrak{g}(Z^{\prime})|\lesssim\frac{r^{4}}{n},$
where the inequality holds up to a constant that depends only on $d_{x}$ and $d_{y}$. Let $\tilde{\mu}_{n}^{\prime}$ and $\tilde{\nu}_{n}^{\prime}$ be the centered empirical distributions corresponding to $Z^{\prime}$. Observe that
$\begin{split}&|\mathfrak{f}(Z)-\mathfrak{f}(Z^{\prime})|\leq|S_{1}(\tilde{\mu}_{n},\tilde{\nu}_{n})-S_{1}(\tilde{\mu}_{n}^{\prime},\tilde{\nu}_{n}^{\prime})|=:\bar{\mathfrak{f}}(Z,Z^{\prime})\quad\text{and}\\ &|\mathfrak{g}(Z)-\mathfrak{g}(Z^{\prime})|\leq|S_{2}(\tilde{\mu}_{n},\tilde{\nu}_{n})-S_{2}(\tilde{\mu}_{n}^{\prime},\tilde{\nu}_{n}^{\prime})|=:\bar{\mathfrak{g}}(Z,Z^{\prime}),\end{split}$
and we will verify that
$\bar{\mathfrak{f}}(Z,Z^{\prime})\vee\bar{\mathfrak{g}}(Z,Z^{\prime})\lesssim\frac{r^{4}}{n}$
up to a constant that depends only on $d_{x}$ and $d_{y}$. Since $\bar{\mathfrak{f}}(Z,Z^{\prime})=r^{4}\bar{\mathfrak{f}}(Z/r,Z^{\prime}/r)$ and $\bar{\mathfrak{g}}(Z,Z^{\prime})=r^{4}\bar{\mathfrak{g}}(Z/r,Z^{\prime}/r)$, it suffices to verify the claim with $r=1$. For $\bar{\mathfrak{f}}(Z,Z^{\prime})$, the desired inequality follows from the decomposition (22) and straightforward calculations. For $\bar{\mathfrak{g}}(Z,Z^{\prime})$, since $(\tilde{\mu}_{n},\tilde{\mu}_{n}^{\prime})$ and $(\tilde{\nu}_{n},\tilde{\nu}_{n}^{\prime})$ are supported in $B_{\mathcal{X}}(0,2)$ and $B_{\mathcal{Y}}(0,2)$, respectively, Lemma 2.2 yields
$\bar{\mathfrak{g}}(Z,Z^{\prime})\leq\sup_{\|A\|_{\operatorname{op}}\leq 1}\big{|}T_{c_{A}}(\tilde{\mu}_{n},\tilde{\nu}_{n})-T_{c_{A}}(\tilde{\mu}_{n}^{\prime},\tilde{\nu}_{n}^{\prime})\big{|}.$
By duality and Lemma 5.4 in [ZGMS24], the right-hand side is $\lesssim n^{-1}$ up to a constant that depends only on $d_{x}$ and $d_{y}$ (see the proof of Proposition 20 in [WB19] for a similar argument).
Now, applying Lemma 3.1, we have
$\mathbb{P}\left(\Delta_{n}\geq 2\big{(}\mathbb{E}[\mathfrak{f}(Z)]+\mathbb{E}[\mathfrak{g}(Z)]\big{)}+Kr^{4}(tn^{-1/2}+\mathfrak{p}_{r})\right)\leq 2\mathfrak{p}_{r}+2e^{-t^{2}},\quad t>0,$
where $K$ is a constant that depends only on $d_{x}$ and $d_{y}$. If $\mu$ and $\nu$ are supported in $B_{\mathcal{X}}(0,r)$ and $B_{\mathcal{Y}}(0,r)$, respectively, then $\mathfrak{p}_{r}=0$ and $\mathbb{E}[\mathfrak{f}(Z)]+\mathbb{E}[\mathfrak{g}(Z)]\lesssim r^{4}\varphi_{n}$ up to a constant that depends only on $d_{x}$ and $d_{y}$ (this follows from Theorem 4.2 in [ZGMS24] or the proof of Theorem 3.1). This establishes Case (i). The rest of the proof is devoted to establishing Cases (ii) and (iii). In what follows, the notation $\lesssim$ means that an inequality holds up to a constant that depends only on $d_{x},d_{y},\kappa,\beta$ and $M$ in Case (ii), and on $d_{x},d_{y},q$ and $M$ in Case (iii). In addition, $K$ denotes a generic constant that depends only on $d_{x},d_{y},\kappa,\beta$ and $M$ in Case (ii), and on $d_{x},d_{y},q$ and $M$ in Case (iii).
Case (ii). In this case, $\mathbb{E}[\mathfrak{f}(Z)]+\mathbb{E}[\mathfrak{g}(Z)]\lesssim\varphi_{n}+n^{-1/2}\sqrt{\log n}$ from the proof of Theorem 3.1 and by taking $q$ large enough and noting that $\mathfrak{m}_{4q}(\mu)\vee\mathfrak{m}_{4q}(\nu)\lesssim 1$. The probability $\mathfrak{p}_{r}$ can be bounded as
$\mathfrak{p}_{r}\lesssim ne^{-(r/M)^{\beta}}.$
Choosing $r=K((\log n)^{1/\beta}+s^{1/4})$ for a large enough $K$ ensures $\mathfrak{p}_{r}\leq(\frac{1}{2})\wedge e^{-s^{\beta/4}}n^{-\kappa}$ and $r^{4}\mathfrak{p}_{r}\lesssim((\log n)^{4/\beta}+s)e^{-s^{\beta/4}}n^{-\kappa}\lesssim n^{-1/2}$ (recall that $s,\kappa\geq 1$).
Case (iii). In this case, $\mathbb{E}[\mathfrak{f}(Z)]+\mathbb{E}[\mathfrak{g}(Z)]\lesssim\bar{\varphi}_{n,q}$ from the proof of Theorem 3.1. The probability $\mathfrak{p}_{r}$ can be bounded as
$\mathfrak{p}_{r}\lesssim nr^{-4q}.$
Choosing $r=Ks^{1/4}n^{1/(4q)}$ for a large enough $K$ ensures $\mathfrak{p}_{r}\leq(\frac{1}{2})\wedge s^{-q}$. This gives the result, finishing the proof. ∎
@@ Appendix A Technical tools
Let $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ be a nonnegative lower semicontinuous function such that there exist nonnegative measurable functions $c_{\mathcal{X}}:\mathcal{X}\to\mathbb{R}_{+}$ and $c_{\mathcal{Y}}:\mathcal{Y}\to\mathbb{R}_{+}$ with $c\leq c_{\mathcal{X}}\oplus c_{\mathcal{Y}}$. Set $\mathcal{B}_{\mathcal{X}}(r)=c_{\mathcal{X}}^{-1}([0,r])$ and $\mathcal{B}_{\mathcal{Y}}(r)=c_{\mathcal{Y}}^{-1}([0,r])$ for $r>0$. Suppose that there exist constants $\kappa>0,\alpha\in(0,1/2]$ and $\delta\geq 0$ such that
$\sup_{(\mu,\nu)\in\mathcal{P}(\mathcal{B}_{\mathcal{X}}(r))\times\mathcal{P}(\mathcal{B}_{\mathcal{Y}}(r))}\mathbb{E}\left[\big{|}T_{c}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c}(\mu,\nu)\big{|}\right]\\ \leq\kappa r(n\wedge m)^{-\alpha}\big{(}\log_{+}(n\wedge m)\big{)}^{\delta}$ (30)
@@
$\sup_{(\mu,\nu)\in\mathcal{Q}_{\epsilon,M}}\mathbb{E}\left[\big{|}T_{c}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c}(\mu,\nu)\big{|}\right]\lesssim(n\wedge m)^{-\alpha}\big{(}\log_{+}(n\wedge m)\big{)}^{\delta}.$
The hidden constant depends only on $\kappa,\alpha,\delta,\epsilon$ and $M$.
@@
We follow the notation used in the proof of Theorem 1.1 in [SH25]. In our case, we may take $s=2$, which yields $\beta=\gamma=1/2$. As such, $\alpha\leq\beta$, so the conclusion follows from Comment 1 after the proof of Theorem 1.1 in [SH25]. The dependence of the constant on the parameters is deduced from the proof. ∎